# Towards a general-purpose foundation model for fMRI analysis

### v1

@@ 3.3 Disease Diagnosis
Figure 4: Evaluation of NeuroSTORM’s performance on disease diagnosis task. (a) Comparison across multiple datasets shows that NeuroSTORM consistently outperforms baseline methods in accuracy, highlighting its strong generalizability for neurological and psychiatric diagnostics. (b) Data efficiency analysis demonstrates that NeuroSTORM maintains robust performance even when trained with only a fraction of the fine-tuning data, underscoring its suitability for data-scarce scenarios.
We evaluated NeuroSTORM’s performance against state-of-the-art methods on the Disease Diagnosis task across multiple datasets, representing different neurological and psychiatric conditions. These datasets include HCP-EP [28], ABIDE [31], ADHD200 [32], COBRE [33], UCLA [34] and motor neuron disease (MND) [30]. As shown in Fig. 4a, NeuroSTORM consistently outperformed all baseline methods including BrainGNN [37], BNT [38], LG-GNN [39], and SwiFT [24] across datasets. On HCP-EP (Control vs Early Psychosis), our model achieved superior accuracy (60.56%) compared to SwiFT’s 58.44% (+2.12%). Similar advantages emerged in ADHD200 (Healthy vs ADHD) where NeuroSTORM attained 60.35% accuracy versus SwiFT’s 57.73% (+2.62%). The performance gap widened substantially in complex multi-class datasets: NeuroSTORM achieved 55.20% accuracy on COBRE (4-class) versus LG-GNN’s 51.44% (+3.76%), and 38.47% on UCLA (4-class) versus SwiFT’s 36.84% (+1.63%). Additionally, on the MND dataset, NeuroSTORM achieved an accuracy of 87.56%, outperforming SwiFT’s 84.83% (+2.73%). These findings indicate that NeuroSTORM demonstrates robust diagnostic capabilities across a variety of neurological conditions when compared to both ROI-based and volume-based approaches.
For data-scarce scenarios, we systematically evaluated NeuroSTORM’s sample efficiency using progressively reduced training subsets (Fig. 4b). With only 30% training data, NeuroSTORM maintained 58.67% and 32.65% accuracy on ADHD200 and UCLA respectively, which are comparable to full-data performance of ROI-based methods like BrainGNN [37] (55.02% and 33.24%). However, performance degraded significantly when using 10% data, particularly on smaller datasets like HCP-EP (insufficient data for evaluation) and COBRE (42.31% accuracy). This contrasts with ROI-based methods’ relative robustness in extreme low-data regimes (BrainGNN: 41.23% with 10% COBRE data), suggesting potential benefits of combining 4D volumetric processing with brain network features. Our experiments reveal that while NeuroSTORM significantly outperforms alternatives at standard data levels (more than 30% training data), hybrid approaches may be warranted for extremely scarce data scenarios.
3.4 fMRI Retrieval
The fMRI retrieval task, for example with the NSD (Natural Scenes Dataset) [35], involves using fMRI data collected while subjects view natural images to establish a correspondence between fMRI sequences and images. The task is to retrieve semantically similar natural images given an fMRI query, or to find relevant fMRI sequences given an image query. The evaluation of fMRI retrieval is crucial for assessing the performance of fMRI foundation models, as it reveals the granularity of image-specific information captured in the predicted brain embeddings. In our experiment, we followed the experimental settings established by Mind-Reader [40], MindEye [4], and MindEyeV2 [41] to evaluate the retrieval performance of fMRI analysis models. As in previous studies, we leveraged a vast pool of image candidates from the LAION-5B [36] dataset, enabling the evaluation to encompass a diverse and fine-grained range of semantics. We began by randomly selecting 300 test samples from NSD dataset and utilized their CLIP embeddings to query the LAION-5B database, storing the top 16 results for each query. Next, we extracted fMRI embeddings for these samples and computed the similarity between the 300 fMRI embeddings and the 4,800 (300 samples $\times$ 16 results) LAION-5B CLIP embeddings to derive retrieval outcomes. Previous works [42, 43, 44, 45, 46, 47, 48, 49, 41, 4, 50] define these results as brain retrieval. By swapping the image CLIP embeddings and fMRI embeddings, we obtained the image retrieval results. We conducted the retrieval process multiple times to ensure robustness and considered both image retrieval and brain retrieval methods. We compared NeuroSTORM with the state-of-the-art volume-based method SwiFT and previous brain decoding methods, including Brain-Diffuser, MindEye, and MindEyeV2. The latter are ROI-based methods that use selected vision-related ROIs to convert 4D volumes into 2D ROI data.
Figure 5: Comparison of NeuroSTORM to other methods on fMRI and image retrieval tasks using the NSD dataset.
Fig. 5 summarizes the results of the fMRI retrieval experiment, comparing the performance of different methods on the NSD [35] dataset for both image and brain retrieval tasks. The results indicate that MindEyeV2 achieved the highest performance in both image and brain retrieval tasks, with accuracies of 98.8% and 98.3%, respectively. MindEye also performed well, achieving an accuracy of 90.0% for image retrieval and 84.1% for brain retrieval. Among the volume-based methods, NeuroSTORM demonstrated notable performance, achieving 73.2% accuracy for image retrieval and 68.3% for brain retrieval, outperforming SwiFT [24], which achieved accuracies of 49.2% and 33.7% for image and brain retrieval, respectively. Brain-Diffuser, an ROI-based method, showed the lowest performance, with 18.8% accuracy for image retrieval and 26.3% for brain retrieval. It was also noted that MindEye and MindEyeV2 utilized only vision-related ROIs, incorporating more biological priors compared to other approaches.
3.5 Task-based fMRI State Classification
Figure 6: Evaluation of Performance on the HCP-YA Dataset for tfMRI state classification task. (a) State classification results: NeuroSTORM demonstrated superior accuracy across all tasks compared to other ROI-based and volume-based methods, achieving an overall accuracy of 92.64%. Task-specific accuracies included 93.56% for Emotion, 91.45% for Gambling, and 94.33% for Language. (b) Performance with varying percentages of fine-tuning data: NeuroSTORM maintained high accuracy even with reduced training data, achieving an overall accuracy of 92.13% with 50% of the data and 88.21% with just 10%. These results highlight its robustness and efficiency in low-resource settings. The variance in the result images was computed based on three independent experimental replicates to ensure statistical reliability.
@@ 3.5 Task-based fMRI State Classification
In this study, we evaluated the tfMRI state classification performance of NeuroSTORM on the HCP-YA [28] dataset. The primary objective of the task is to accurately identify the cognitive state or experimental condition. The HCP-YA dataset comprises seven distinct functional tasks: Emotion, Gambling, Language, Motor, Relational, Social, and Working Memory (WM). Each task is designed to activate different brain networks associated with unique cognitive states. NeuroSTORM exhibited superior performance across all tasks (Fig. 6), highlighting its ability to effectively generalize features. Our framework achieved an overall accuracy of 92.64%, significantly outperforming both ROI-based approaches and recent volumetric deep learning methods. Specifically, NeuroSTORM surpassed the previous state-of-the-art volumetric method BAnD [51] (91.74%) by 0.9% accuracy, while showing 23.4% and 11.4% improvements over traditional BrainGNN [37] (69.73%) and BNT [38] (81.22%) approaches respectively. Task-specific comparisons further evaluated the model capacity in complex cognition tasks. NeuroSTORM achieved 94.33% accuracy in Language classification (+0.46% over BAnD) and 93.56% in Emotion recognition (+1.78% over BAnD).
Additionally, we investigated NeuroSTORM’s performance robustness in low-resource scenarios. Using only 50% of training data, NeuroSTORM maintained 92.13% overall accuracy, significantly exceeding BAnD’s full-data accuracy. Even with 10% training data, NeuroSTORM achieved 88.21% accuracy, wchih is higher than BrainGNN’s 69.73% with full data. Notably, task-specific performance remained stable across data reductions, with Gambling task accuracy decreasing only 2.89% (from 91.45% to 88.56%) at 10% training data. These results underscore the robustness and efficiency of NeuroSTORM in low-resource settings.
@@ 4 Discussion
Our study addresses two fundamental challenges in fMRI research, the limited model reproducibility and transferability, through systematic innovations in data curation, architecture design, and task benchmarking. The construction of the largest multi-source fMRI pre-training corpus to date ($>65,000$ scans across 15 sites) combined with the STRD module and SWM architecture establishes NeuroSTORM as a versatile foundation model. NeuroSTORM directly processes raw 4D volumes while mitigating the influence of scanner artifacts and physiological noise. By formulating fMRI feature learning as spatiotemporal reconstruction invariant to population-level heterogeneity, we demonstrate that large-scale pre-training can simultaneously mitigate fundamental issues in test-retest reliability and cross-task generalization. The comprehensive benchmark spanning five domains (sociodemographic prediction, phenotype estimation, disease diagnostics, content retrieval, and state classification) systematically evaluates general-purpose fMRI modeling capabilities, covering clinical, neuroscientific, and BCI applications.
@@
Unlike ROI-based approaches that discard anatomical context through atlas parcellation [5, 6], NeuroSTORM’s volume-level processing preserves spatiotemporal relationships critical for modeling brain activities. The STRD module advances noise resilience by reconstructing both brain signals and noise from heterogenous population during masked pre-training. Moreover, we design a fMRI analysis benchmark which meets key requirements for evaluating foundation models in neuroimaging: (1) task diversity (regression/classification across neurodevelopmental, psychiatric, and neurodegenerative conditions), (2) biological plausibility (alignment between predicted phenotypes and known neuromarkers), and (3) clinical transferability (efficacy in low-data regimes through prompt tuning). The inclusion of retrieval and tfMRI classification tasks further tests the model’s capacity to decode fine-grained semantic content, which is absent in the benchmark of previous fMRI analytical models.
@@
NeuroSTORM demonstrates statistically superior performance across all benchmarking tasks compared to state-of-the-art methods. For disease diagnosis, it achieves 75.2% accuracy on schizophrenia detection (HCP-EP) and 60.4% AUROC differentiating ADHD subtypes (UCLA), outperforming ROI-based approaches by 15.3% and 22.7% respectively. The model attains 93.3% gender classification accuracy (HCP-YA) with 3.4 years MAE in age prediction, reducing errors by 38% versus SwiFT [24] through spatiotemporal redundancy dropout. Phenotype prediction reveals strong alignment with established biomarkers, such as the 0.429 PCC for emotion task. Notably, NeuroSTORM achieves 79.4% brain-to-image retrieval accuracy (NSD dataset), approaching specialized vision-language models like MindEyeV2 (98.3%) while using raw volumes rather than annotated ROIs. Task fMRI classification reaches 92.6% accuracy across seven different states, confirming effective adaptation to both resting-state and task activity patterns. Notably, we observe that ROI-based methods exhibit lower performance compared to volume-based approaches, due to the loss of state-related information during downsampling operations; this finding is also consistent with reported results in previous works [51].
NeuroSTORM’s parameter-efficient prompt tuning requires $\leq 3$% tuned weights while maintaining efficient data utilization. For instance, NeuroSTORM achieves 92.1% state classification accuracy with 50% training data, which is a critical advantage in data-scarce scenarios. This efficiency stems from pre-training’s implicit denoising: visualization of attention maps reveals suppressed physiological artifacts in motion-corrupted scans, suggesting intrinsic noise cancellation. Clinically, the 60.6% accuracy distinguishing schizophrenia subtypes (COBRE dataset) with 30 training samples approaches human inter-rater reliability, demonstrating potential for automated differential diagnosis.

### v2

@@ 3.3 Disease Diagnosis
We evaluated NeuroSTORM’s performance against state-of-the-art methods on the Disease Diagnosis task across multiple datasets, representing different neurological and psychiatric conditions. These datasets include HCP-EP [28], ADHD200 [31], COBRE [32], UCLA [33] and motor neuron disease (MND) [30]. As shown in Fig. 4a, NeuroSTORM consistently outperforms all baseline methods across datasets. On ADHD200 (healthy vs ADHD) it achieves 71.46%, 1.38% higher than the best competitor BrainLM. For the COBRE benchmark NeuroSTORM reaches 69.84%, surpassing Com-brainTF by 2.44%. In early-psychosis detection (HCP-EP) the model obtains 69.06%, a 1.44% gain over Com-brainTF. On the MND cohort NeuroSTORM scores 86.50%, improving on SwiFT by 1.74%. The largest margin was observed on the UCLA dataset, where NeuroSTORM attained 50.66%, 2.62% above TFF and more than 4% ahead of all volume baselines. These results demonstrate NeuroSTORM’s robust and generalizable diagnostic capability across diverse brain disorders [41, 42].
For label efficiency, we systematically evaluated NeuroSTORM’s sample efficiency using progressively reduced training subsets (Fig. 4b). On ADHD200 the model retains 68.7% accuracy with only 50% of the labels, already surpassing Com-brainTF trained on all data (68.0%) and nearly matching the best volume baseline BrainLM (70.1%). On the COBRE dataset NeuroSTORM delivers 62.4% accuracy with 30% of the labels, only 2.1% below SwiFT’s full-data score of 64.5% and close to BrainLM’s 64.1%. These examples illustrate that NeuroSTORM’s performance degrades more gradually as labelled data decrease, and that it maintains comparable accuracy to state-of-the-art methods in data-scarce scenarios.
3.4 fMRI Re-identification
To evaluate the discriminative strength of NeuroSTORM’s embedding space, we follow a closed-set fMRI re-identification protocol in which the held-out data are split into a query set and a non-overlapping gallery set, with the gallery size fixed at 100. During inference, each 4D sequence is forwarded once to obtain a single $\ell_{2}$-normalized feature vector, and an exhaustive nearest-neighbor search is performed in the gallery. Retrieval quality is reported with the standard Rank-1 Accuracy and mean Average Precision (mAP) over the entire query set. This experimental design offers a rigorous test of how well the foundation model retains subject-specific cues [43] and supports large-scale similarity search in high-dimensional fMRI data.
Figure 5: Evaluation of fMRI re-identification on HCP-YA with the gallery size fixed at 100. Under a closed set protocol, each 4D sequence is embedded once into an $\ell_{2}$-normalized feature vector and matched via exhaustive nearest neighbor search in the gallery, with retrieval quality reported by Rank-1 Accuracy and mAP.
As shown in Fig. 5, NeuroSTORM achieves 93.1% Rank-1 accuracy and 92.4% mAP under the same closed-set protocol, surpassing all competing methods. Compared with the strongest baseline SwiFT, the gains are +1.5% in Rank-1 and +3.0% in mAP. ROI-based baselines show lower performance. For example, the best of ROI-based model, BrainLM, reaches 90.8% Rank-1 and 88.2% mAP, whereas Com-brainTF attains 89.0% and 87.4%. The other methods obtain even lower scores. These margins suggest that the compact $\ell_{2}$-normalized embeddings produced by our shifted scanning strategy on a Mamba backbone preserve strong subject-specific signatures. The spatial-temporal optimized pre-training coupled with task-specific prompt tuning further supports reliable one-to-many retrieval in high-dimensional fMRI space.
3.5 fMRI State Classification
@@
In this study, we evaluated the state classification performance of NeuroSTORM in the HCP-YA [28] and DMT-HAR-MED [34] dataset. The primary objective of the task is to accurately identify the cognitive state or experimental condition. The HCP-YA dataset comprises seven distinct functional tasks: Emotion, Gambling, Language, Motor, Relational, Social, and Working Memory (WM). Each task is designed to activate different brain networks associated with unique cognitive states. The DMT-HAR-MED dataset includes resting-state fMRI from 40 meditators collected before and after a retreat, with participants receiving either DMT-harmine or placebo.
Figure 6: Evaluation of Performance for fMRI state classification task. (a) tfMRI state classification results: NeuroSTORM demonstrated superior accuracy across all tasks compared to other ROI-based and volume-based methods, achieving an overall accuracy of 92.64%. (b) Medication state classification results: NeuroSTORM distinguished DMT-harmine from placebo with 87.5% overall accuracy. The violin plots display the distribution of predicted probabilities assigned to the true class for each subject; correct predictions cluster at high-confidence values while errors occur at markedly lower probabilities, indicating well-calibrated predictions. Variances were estimated from five technical replicates. Pairwise significance markers were computed using a two-sided paired t-test (ttest_rel) without multiple-comparison correction, and the corresponding P-value is annotated.
NeuroSTORM exhibited superior performance across all task states (Fig. 6(a)), highlighting its ability to effectively generalize features. Our framework achieved an overall accuracy of 92.40%, outperforming the strongest volume-based baseline BAnD (91.46%) and showing clear margins over the best ROI baselines BNT (80.62%). Task-specific comparisons further evaluated the model capacity in complex cognition tasks. NeuroSTORM achieved 93.90% accuracy in Language (+0.78% over BAnD) and 93.06% on Emotion (+1.56%). Additionally, we investigated NeuroSTORM’s performance robustness in data-scared scenarios. NeuroSTORM retained 91.2% overall accuracy with 50% of the labels, matching BAnD’s full-data score of 91.5% and remaining far ahead of the strongest ROI baseline BNT at full data (80.6%). Even with only 10% of the labels NeuroSTORM still reached 84.2%, whereas BAnD fell to 77.4% and BNT to 67.0%. These results underscore the label efficiency of NeuroSTORM in data-scared settings.
In the DMT-HAR-MED [34] dataset, we evaluated NeuroSTORM for medication state classification between DMT harmine and placebo. As shown in Fig. 6(b), NeuroSTORM achieved 90.0% average accuracy in five independent experiments, demonstrating effective detection of medication related functional changes in resting state. Confidence analysis shows that correct predictions have probability outputs concentrated at high values, whereas incorrect predictions have markedly lower confidence. These results confirm that NeuroSTORM delivers high accuracy together with high confidence and stable predictive behavior in medication state classification, highlighting robustness and practical utility in resting state drug intervention settings.
@@ 4 Discussion
Our study addresses two fundamental challenges in fMRI research, the limited model reproducibility and transferability, through systematic innovations in data curation, architecture design, and task benchmarking. The construction of the large multi-source fMRI pre-training corpus (>50,000>50,000 scans) combined with the STRD module and SWM architecture establishes NeuroSTORM as a versatile foundation model. NeuroSTORM directly processes raw 4D volumes while mitigating the influence of scanner artifacts and physiological noise. By formulating fMRI feature learning as spatiotemporal reconstruction invariant to population-level heterogeneity, we demonstrate that large-scale pre-training can simultaneously mitigate fundamental issues in test-retest reliability and cross-task generalization. The comprehensive benchmark spanning five domains (sociodemographic prediction, phenotype estimation, disease diagnostics, fMRI re-identification, and state classification) systematically evaluates general-purpose fMRI modeling capabilities, covering clinical, neuroscientific, and BCI applications.
Unlike ROI-based approaches that discard anatomical context through atlas parcellation [5, 6], NeuroSTORM operates directly on whole-brain 4D fMRI volumes, thereby preserving the fine-grained spatiotemporal relationships that are critical for modelling neural activity. This volume-level strategy need substantial computational efficiency, so we adopt a Shifted-Window Mamba backbone that can process high-dimensional inputs within practical GPU memory limits. During downstream adaptation we further reduce the computational footprint through task-prompted tuning, which updates fewer than 5% of the parameters while maintaining full model capacity. A second technical challenge is the intrinsic redundancy of fMRI signals. The STRD module addresses this issue by selectively masking spatially and temporally redundant voxels during masked pre-training and then reconstructing both informative brain signals and subject-specific physiological noise. These combined innovations enable NeuroSTORM to learn long-range spatiotemporal dependencies and to produce a transferable feature space that generalises across tasks, datasets and computational budgets. Consequently, NeuroSTORM provides a flexible foundation for future fMRI research and downstream modelling efforts under diverse real-world conditions.
@@
Moreover, we design a fMRI analysis benchmark which meets key requirements for evaluating foundation models in neuroimaging: (1) task diversity (regression/classification across neurodevelopmental, psychiatric, and neurodegenerative conditions), (2) feature-space validity (demonstrated by the structured embeddings uncovered in the fMRI re-ID task), and (3) clinical transferability (efficacy in low-data regimes through prompt tuning). The inclusion of re-identification and state classification tasks further tests the model’s capacity to decode fine-grained semantic content, which is absent in the benchmark of previous fMRI analytical models.
@@
NeuroSTORM demonstrates statistically superior performance across all benchmarking tasks compared to state-of-the-art methods. For disease diagnosis, it achieves 69.06% accuracy on schizophrenia detection (HCP-EP) and 50.66% accuracy differentiating ADHD subtypes (UCLA), outperforming ROI-based approaches by 1.44% and 3.54% respectively. The model attains 93.28% gender classification accuracy (HCP-YA) with 8.24 MAE in age prediction, reducing errors by 5.5% versus SwiFT [24]. Phenotype prediction reveals strong alignment with established biomarkers, such as the 0.419 PCC for emotion task in HCP-YA dataset. fMRI state classification reaches 92.4% accuracy across seven different states, confirming effective adaptation to both resting-state and task activity patterns. Notably, we observe that ROI-based methods exhibit lower performance compared to volume-based approaches, due to the loss of state-related information during downsampling operations; this finding is also consistent with reported results in previous works [44]. Furthermore, NeuroSTORM surpasses the transformer-based SwiFT on the vast majority of tasks, this comparison validates the effectiveness of our SWM backbone; Sec. 5 further shows that NeuroSTORM attains these gains while consuming less GPU memory, underscoring the computational efficiency of the proposed architecture. This improvement benefits from the window mixing and state space model design that captures long-range spatiotemporal dependencies without incurring the quadratic complexity of full self-attention.
NeuroSTORM’s parameter-efficient prompt tuning requires $\leq 5\%$ tuned weights while maintaining efficient data utilization. For instance, NeuroSTORM achieves $91.18\%$ state classification accuracy with $50\%$ of the training data in HCP-YA dataset, which is a critical advantage in data-scarce scenarios. When the available data are further reduced to $30\%$, the model still attains $88.04\%$ overall accuracy, only $4.36$ percentage points below the $92.40\%$ obtained with the full dataset. Even with merely $10\%$ of the data, NeuroSTORM retains $84.24\%$ overall accuracy, and every individual task remains above $83\%$. These findings confirm that the proposed prompt-tuning scheme preserves balanced generalization across cognitive states and degrades gracefully under severe data-limitation settings.